OpenStudy (immings):

Solve for X : not looking for an Answer need to know how to approach the problem. I have an idea, but unsure if correct. 2x^3+6x^2-x=3

OpenStudy (anonymous):

factor out x

OpenStudy (ajprincess):

2x^3+6x^2-x=3 Bring all the terms to one side 2x^3+6x^2-x-3=0 Nw try to factorise 2x^3+6x^2 and -x-3. Can u tell me vat u get @immings?

OpenStudy (immings):

2(x^3+3x^2) -(x+3)

OpenStudy (ajprincess):

2(x^3+3x^2) -(x+3) U can factor out x^2 from x^3+3x^2.

OpenStudy (immings):

ahh, forgot to do that. 2x^2(x+3) -(x+3)

OpenStudy (immings):

so I can cancel out (x+3) and multiply (2x^2)(-1)?

OpenStudy (ajprincess):

No u can't cancel it. U hav to factorise it. To make it simple let us assume u=x+3. 2x^2(x+3) -(x+3) 2x^2u-u Nw can u factorise it?

OpenStudy (immings):

I don't see what I can factor out here.

OpenStudy (immings):

an x maybe?

OpenStudy (ajprincess):

u can factor out u from 2x^2u-u. Then u get u(2x^2-1) u=(x+3) so (x+3)(2x^2-1). clear?

OpenStudy (immings):

Yes. it makes more sense this way

OpenStudy (ajprincess):

Nw (x+3)(2x^2-1)=0 Either x+3=0 or (2x^2-1)=0 Can u find x nw?

OpenStudy (immings):

Yes. Thank you. when I got to this step I was using LD to try to solve it. :s

OpenStudy (ajprincess):

oh k. yw:)